### v1

# RefExpo: Unveiling Software Project Structures through Advanced Dependency Graph Extraction

Abstract.
Assessing the dependency graph (DG) of a software project offers valuable insights for identifying its key components. Numerous studies have explored extracting DGs and leveraging them for various analyses, including security and bus factor calculations. However, there is a lack of user-friendly tools for DG extraction, and no comprehensive DG datasets from open-source projects are available.
This study introduces RefExpo, an easy-to-use DG extraction tool supporting multiple languages like Java, Python, and JavaScript. Based on the IntelliJ plugin SDK, RefExpo ensures compatibility with various project structures and technology versions. We also provide a dataset of 20 Java and Python projects, with plans to expand upon request.
To validate RefExpo  we focused on Java and Python. Our tests showed RefExpo achieving 92% and 100% recall on micro test suites Judge and PyCG for Python and Java, respectively. In macro-level experiments, RefExpo outperformed existing tools by at least 31% and 7% in finding unique and shared results. You can access the source code of our tool from our replication package1. The installable version of RefExpo is available on the IntelliJ marketplace3. Additionally, a short video describing its functionality can be viewed here: https://youtu.be/eCnPUlj6YgA.
Software Analytics, Dependency Graph, Call Graph, Code Referencing, Software Structure, Software Architecture
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 39th IEEE/ACM International Conference on Automated Software Engineering; Oct 27– Nov 1, 2018; Sacramento, CAisbn: 978-1-4503-XXXX-X/18/06booktitle: Companion Proceedings of the 39th IEEE/ACM International Conference on Automated Software Engineering (ASE 2024)$c$s: Software and its engineering Software system structuresccs: Software and its engineering Software architecturesccs: Software and its engineering Abstraction, modeling and modularityccs: Software and its engineering Compilers
1. Introduction
In the recent years, numerous studies have focused on analyzing the dependency graph (DG) of software projects (Bhattacharya et al., 2012; Ma et al., 2018; Deng and Jones, 2012; Cai and Santelices, 2015; Chen et al., 1997; Walunj et al., 2019; Ryder, 1979; Salis et al., 2021; Reif et al., 2016; Keshani et al., 2023). A dependency graph typically represents software components as nodes and their logical dependencies as edges. In such a graph, nodes can be methods, classes, files, modules, etc. Logical dependencies are identified by how each component references others. These references can be method calls, property access, inheritance relationships, etc.
Meanwhile, there are other studies that primarily aim at utilizing DG to extract insights about the significance (Inoue et al., 2005; Wang et al., 2022; Pan et al., 2021; Li et al., 2021), comprehensibility (Şora, 2016; Kamran et al., 2013; Yang et al., 2016; Thung et al., 2014), and influence of software project components (Feng and Maletic, 2006; Li et al., 2018; Gu et al., 2014). Moreover, many of the modern software analytics tools such as Sonargraph (son, [n. d.]) and numerous studies integrate dependency data into their analytical frameworks, utilizing it to focus their assessments on the most critical components of the project (Zimmermann and Nagappan, 2007, 2008; Nguyen and Tran, 2010; Du et al., 2022; Meilong et al., 2020). Despite the abundance of such studies, there is a notable lack of user-friendly tools for generating dependency graphs (Salis et al., 2021). This gap highlights the technical challenges researchers face in extracting such data, often restricting them to simpler projects in specific technologies like Java.
There are many tools, both open-source and commercial, for creating dependency graphs. However, they are often difficult to use. Some of these tools are outdated and do not work with the latest versions of the languages they support.(pya, [n. d.]; Salis et al., 2021; pyn, [n. d.]) For example, we observed that the majority of the Python tools such as PyCG (Salis et al., 2021) and Pyan (pya, [n. d.]) are not compatible with the current version of Python (3.12 at the time). These tools could not be run on actively maintained popular GitHub repositories. Nevertheless, we tried to use older versions of the repositories but it seems that the tools are highly sensitive to the Python version and we only could find one revision in an experimental repository.
Besides, some of the tools function adequately only for small projects and are difficult to configure for larger ones. (jar, [n. d.]; son, [n. d.]; df, [n. d.]; nde, [n. d.]) Although the definition of a small-sized project can vary, we have found it challenging to use these tools on some popular GitHub repositories. For example, a big project like ElasticSearch does not compile into a simple WAR file that these tools can process. Besides, some code that can be executed during the build process, such as test cases, can be a part of the dependency graph. In addition, the accuracy of some of the existing tools might be inadequate for some of the use cases. (rex, [n. d.])
Furthermore, configuring tools that provide evaluation over source code can be difficult. For instance, Sonargraph (son, [n. d.]) requires manual configuration over the build tools to be able to provide its evaluations. This can be done easily on small projects; however, for a larger project with multiple levels of submodules doing so is not an easy task. For instance, the complex build structure of ElasticSearch makes it difficult to configure Sonargraph probes, and we were unable to do so in our attempts. Besides, we observed that the provided tools for Python such as PyCG (Salis et al., 2021) and Pyan (pya, [n. d.]) are not compatible with the current version of Python (3.12 at the time). We could not run any of them over popular GitHub repositories since they are actively maintained. Nevertheless, we tried to use older versions of the repositories but it seems that the tools are highly sensitive to the Python version, and we only could find one revision in an experimental repository.
In this work, to facilitate dependency extraction, we present an easy-to-use tool called RefExpo  [^1] https://github.com/vharatian/RefExpo, which has been utilized in our previous study BFSig (Haratian et al., 2023). RefExpo is based on IntelliJ IDEA, one of the most robust Integrated Development Environments (IDEs) available. IDEs like IntelliJ are highly effective at evaluating dependency graphs because they offer robust code navigation and reference identification. Since IDEs are designed for compatibility, they stay updated with the latest language versions. Besides, their user-friendly interfaces make importing and managing projects straightforward and much easier than the experimental tools.
In addition, the unavailability of reliable datasets might compromise the validity of studies, as the technical nature of data extraction is sensitive to minor changes that can significantly impact study outcomes (Du et al., 2022). This challenge inspired us to create a dataset to potentially accelerate research in this domain. This allows researchers who want to include DGs in their projects to concentrate on their studies without the burdens of DG extraction. We have compiled an initial dataset1 of 20 popular GitHub projects, encompassing Java, and Python, with an eye toward expansion.
2. RefExpo
RefExpo generates comprehensive dependency graphs from projects developed in a variety of programming languages. Utilizing InteliJ as the underlying engine, RefExpo can support projects consisting of multiple programming languages including Java, Python, JavaScript, Go, Ruby, and HTML, and also framework-specific syntax. For instance, it is possible to add Python support by adding a Python plugin, which is provided by the internal JetBrains team. In addition, by adding framework support (for instance, Angular), RefExpo can provide a more sophisticated dependency graph for custom framework implementations. Finally, the utilization of an advanced engine such as IntelliJ allows RefExpo to evaluate large-scale projects with zero to no required configurations. Nevertheless, in large projects with custom structures, some configurations might be required. However, due to the popularity of IntelliJ, these projects usually share a guide on how to configure the project for IntelliJ.
One difference of RefExpo compared to other available dependency extraction tools is its inability to trace dynamic references between components. Although extracting dynamic references can be crucial for some applications, it is acceptable in most cases. We also demonstrate the accuracy of our tool is sufficient in dynamic programming environments where dynamic referencing is more prevalent. In parts of the literature (Ryder, 1979; Walunj et al., 2019), graphs that include dynamic links are referred to as call graphs, while those containing only static references are called dependency graphs. We will continue using these terms accordingly. A detailed discussion of the differences between these graphs is included in Appendix B [^2] https://figshare.com/articles/journal_contribution/Appendix_Performance_Evalaution_of_RefExpo/26110708, .
The tool comes in the format of a plugin that can be installed on IntelliJ through its marketplace [^3] https://plugins.jetbrains.com/plugin/23684-refexpo, or could be used with IntelliJ headless mode in the command line through the source code1. To use RefExpo, the user needs to import their project into IntelliJ and ensure that IntelliJ is properly configured to index the source code, and then open the RefExpo toolbox. No project build is required, but to make sure that the project is imported correctly and is indexed by IntelliJ, one can try to build the project before running the evaluation. To manage large projects effectively and enhance usability, RefExpo offers configurable filters. The filters allow users to apply regular expressions for evaluating files, classes, and methods. Additionally, RefExpo detects circular references and provides options to filter them out at the file, class, and method levels. While these filters could be applied to the CSV file after the evaluation, applying a filter beforehand could help avoid generating an excessively large output file. Besides, evaluating all files with all references in a large project might result in long execution which can be sped up by prior filtering.
Finally, the tool produces a CSV file that contains all identified references, categorized by three locators: File, Class, and Method, for both the source and the target. You can refer to a complete usage guide including a short video, available in the replication package1.
3. Implementation
To understand the process and intricacies of RefExpoś inspection mechanism, it is essential to delve into its two-phase strategy.
RefExpo has two inspection phases that are quite similar to each other. Due to the lazy indexing notion of IntelliJ, RefExpo needs to administer a two-pass strategy. Lazy indexing in IntelliJ postpones the indexing process until a file is opened and indexing is needed, rather than keeping the user waiting for the entire indexing to complete. This is in contrast to what RefExpo requires, which is the complete indexing before running the evaluations. Therefore, in the first path, RefExpo will traverse all possible files and do a full traverse over their Abstract Syntax Tree (AST), which is provided by the PSI API of IntelliJ plugin SDK (int, [n. d.]). For each element in the AST, IntelliJ provides a list of places in which the statement references. In the first path, RefExpo only resolves all the references, which makes the IntelliJ index target files of the reference and will keep this information. Afterward, RefExpo does another traversing round, which will go over all files and elements again. However, this time, it will evaluate each resolved reference and log its encounter. Figure 1 shows the general structure of RefExpo inspection.
RefExpo focuses on internal referencing and does not evaluate external references. For instance, it is usual to reference java.lang.String in a Java project, however, RefExpo filters out all those references. To achieve this, RefExpo initially filters out project files to those included in the version control system (VCS) of the project. This will also exclude the build products in addition to the dependency files since IntelliJ is capable of reading and providing references in already compiled CLASS and JAR files. Finally, the target of each reference should be included in the VCS so RefExpo consider it as a valid reference in the project.
Ultimately, RefExpo generates a CSV file listing all encountered references. Each reference will include three locators – File, Class, and Method – for both the source and target. Depending on the reference type, it may have one, two, or all three of these locators. For instance, since HTML files do not contain classes or methods, a reference to or from an HTML file will be labeled only by the file name. Consequently, the class and method columns will be empty in the CSV file.
Figure 1. RefExpo general structure
4. Dataset
To help researchers quickly access an extracted dependency graph, we have evaluated 20 repositories developed in Python and Java 1. Our selection process involves identifying popular projects on GitHub and utilizing its search API. We begin by ranking repositories according to their star count and subsequently gather data on their programming languages. Following this, we specifically filter for repositories primarily developed in Java and Python. We manually evaluated and removed repositories that are small code samples from the resulting repositories. Finally, we have added those projects that fall out of this filtering process but we have evaluated them during our macro evaluation phase. Those projects were initially selected due to their compatibility with other tools which we compared RefExpo results with. As it has been presented in our replication package 1, the dataset contains projects with different sizes from only 11 files to 15K. This approach yields a comprehensive dataset designed to assist researchers with diverse requirements. Our objective is to continually expand this dataset, and we are open to collaborations with researchers, offering to tailor the dataset to meet their specific needs.
5. Evaluation
We evaluated RefExpo to assess its accuracy and effectiveness in identifying dependencies by comparing it against other tools. The evaluation was conducted at both micro- and macro levels, using two test suites and eight popular GitHub repositories, respectively.
Micro-level focuses on detailed, small-scale tests to validate specific functionalities and identify any bugs or inaccuracies. we used the Judge and PyCG data suites. Although initially designed for call graphs, we adapted these suites for RefExpo by refining them to focus on relevant cases, removing test cases involving dynamic and external links. Our results indicate that:
Micro Evaluation Results RefExpo achieved 100% recall on the cleaned version of the Judge Test Suite, and 85% on the initial suite. For the PyCG Test Suite, it achieved 97% recall on the cleaned version and 68% on the initial suite.
Macro-level involves large-scale tests on real-world projects to evaluate the overall performance and robustness of RefExpo. We compared RefExpo against other tools over eight popular GitHub repositories to see how well it performs in practical scenarios.
Macro Evaluation Results For Python projects, RefExpo covered 90% of identified edges, outperforming Pyan (33%) and PyCG (26%). RefExpo had 35% shared and 56% unique edges. For Java projects, RefExpo covered 92% of identified edges, surpassing Jarviz (41%), Dependency Finder (29%), and Sonargraph (55%). RefExpo had 59% shared and 34% unique edges.
In both evaluations, RefExpo demonstrated superior recall and accuracy compared to other tools, indicating its effectiveness in accurately identifying dependencies. A detailed description of our evaluation methods, metrics, complete results table, and accompanying diagrams can be found in Appendix C2
6. Conclusion
Dependency graphs contain valuable information that could be incorporated into software analytics studies. However, due to the shortage of an open-source, user-friendly tool capable of generating reliable dependency graphs, dependency graphs are not present in as many studies as they can. Existing tools are often challenging to use and limited in their support for various programming languages and artifacts. RefExpo has been developed to address this shortfall, leveraging the well-documented and community-supported tool IntelliJ. This integration enables RefExpo to support a diverse array of programming languages, ranging from Java to HTML.
Our evaluations suggest that RefExpo achieves 92% and 100% recall on micro test suites Judge and PyCG for Python and Java, respectively. In macro-level experiments, RefExpo outperformed existing tools by at least 31% and 7% in finding unique and shared results. Additionally, we have commenced the creation of a dataset encompassing dependency graphs from 20 popular GitHub repositories, consisting of Java and Python projects.
Acknowledgements
This study was mainly supported by the Scientific and Technological Research Council of Turkey (TUBITAK) 3501 program (Project Number: 121E584) and partially supported by JetBrains Research. Special thanks to JetBrains’ IntelliJ team for their help in maximizing the potential of the tool.